OpenStudy (anonymous):

Can some one please help me solve the following problems i am working on functions. 1)y=2x+7 2)y=x2-7 3)y=x-45+3x 4) (2,4) (6,10) (12,_?_) 5) (7,_?_) (10,12) (12,16)

OpenStudy (anonymous):

1)x=(y-7)/2 2)x=+-√(y+7)/4 3)x=(y+45)/4 4)18 5)10

OpenStudy (anonymous):

thanks for the help means alot (:

OpenStudy (anonymous):

you r always welcome!

OpenStudy (anonymous):

i'm sorry,the second answer is wrong,it is, 2)+-√(y-7) unknowingly,i put a 4 there...sorry...

OpenStudy (anonymous):

OKAY THANKS(: